WACC and Percentage of Debt Financing Hook Industries' capital structure consists solely of debt and common equity. It can issue debt at rd = 11%, and its common stock currently pays a $2.00 dividend per share (D0 = $2.00). The stock's price is currently $28.00, its dividend is expected to grow at a constant rate of 6% per year, its tax rate is 35%, and its WACC is 12.55%. What percentage of the company's capital structure consists of debt? Round your answer to two decimal places. %
Answer #1

Cost of Equity = [{D0 * (1 + g)} / P0] + g

= [{$2 * (1 + 0.06)} / $28] + 0.06 = 0.0757 + 0.06 = 0.1357, or 13.57%

wD + wE = 1

wE = 1 - wD

WACC = [wD * kD * (1 - t)] + [wE * kE]

12.55% = [wD * 11% * (1 - 0.35)] + [(1 - wD) * 13.57%]

12.55% = [wD * 7.15%] + 13.57% - [13.57% * wD]

12.55% - 13.57% = -[wD * 6.42%]

-1.02% = -[wD * 6.42%]

wD = 1.02%/6.42% = 0.1591, or 15.91%